OpenStudy (anonymous):

I have the answer can someone please check me... Find the derivative p(x)=10x^2sin(10x)

OpenStudy (anonymous):

this isi the equation for the question\[p(x)=10x^2\sin(10x)\]

OpenStudy (anonymous):

product plus chain rule

OpenStudy (anonymous):

\((fg)]=f'g+g'f\) with \(f(x)=x^2,f'(x)=2x,g(x)=\sin(10x),g'(x)=10\cos(x)\)

OpenStudy (anonymous):

damn i meant \(g'(x)=10\sin(10x)\)

OpenStudy (anonymous):

and factor that first ten out of the whole thing

OpenStudy (anonymous):

I got 20xsin(10x)+100x^2cos(10x)

OpenStudy (anonymous):

looks good to me

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

if you want to be fancy you can factor \(20x\) out of both terms